OpenStudy (anonymous):

f(x)= 3x^2/ 2(x-3)(x+2) What is the Horizontal asymptote?

OpenStudy (anonymous):

\[f(x) = \frac{3x^2}{2x^2 - 2x - 12}\] So, \[x \to \infty \implies f(x) \to \frac{3}{2}\]